OpenStudy (anonymous):

factor 3x^2+2x-8 step by step?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

3x^2+2x-8=0 ax^2+bx+c=0 you want to look for 2 numbers that multiply to (a times c) and add to b so that would be two numbers that multiply to -24 and add to 2 in this case -4 and 6, these will become your new coefficients of x for b 3x^2 +6x -4x -8 =0 ,now you can factor by grouping 3x(x+2)-4(x+2) =0 (x +2)(3x-4)=0 x= -2 , x= 4/3

OpenStudy (anonymous):

thank you